Given that cos 80 = 0.173648... explain how to find cos100, cos260, cos280, and sin190 without using a calculator.
Answer by ikleyn(52847)

.

cos(100°) = -cos(80°) = -0.173648...


cos(260°) = cos(180°+80°) = -0.173648...


cos(260°) = cos(360°-80°) = cos(80°) = 0.17648...


sin(190°) = cos(190°-90°) = cos(100°) = -cos(80°) = -0.173648...


Answered and solved.